OpenStudy (quickstudent):

Can someone please help me solve this? 5,184 = (6m)^2

OpenStudy (phi):

(6m)^2 means 6*m*6*m

OpenStudy (phi):

your problem is 6*6*m*m= 5184 I would simplify 6*6

OpenStudy (quickstudent):

Oh, ok. I'll try solving starting with that.:)

OpenStudy (quickstudent):

Now I got another problem. I'm not sure how to solve further then this. 5,184 = 6 * 6 * m * m 5,184 = 36m * m 5,184/36 = 36m/36 * m 144 = mm How do I solve this now?

OpenStudy (phi):

so far, so good m*m= 144 or \[ m^2 = 144\] people take the square root of both sides

OpenStudy (quickstudent):

Oh, so the answer will be 12 = m. Right?

OpenStudy (phi):

yes. m=12

OpenStudy (phi):

if you put 12 in for m in the original problem (6m)^2 to get (6*12)^2 = 72^2 = 5184, it works. that shows m=12 is correct.

OpenStudy (quickstudent):

OK, I understand now. Thanks!